OpenStudy (anonymous):

The average annual salary of the employees of a company in the year 2005 was $80,000. It increased by the same factor each year and in 2006, the average annual salary was $88,000. Let f(x) represent the average annual salary, in thousand dollars, after x years since 2005. Which of the following best represents the relationship between x and f(x)?

OpenStudy (anonymous):

f(x) = 88(0.88)x f(x) = 88(1.1)x f(x) = 80(0.88)x f(x) = 80(1.1)x help

OpenStudy (anonymous):

Since x is the number of years after 2005, then x=1 for 2006. So if you just plug in x=1 into the four answers, only the last one will give you 88, for 88,000. So the answer is the last one.